kaylak:

determine whether f(x0=5x+1/x and g(x)-x/5x+1 are inverses with steps y'all

kaylak:

@Falconmaster

Falconmaster:

i dont think so it does not look like it

kaylak:

have any steps i know it isn't but i have to show work

Falconmaster:

oh umm mathway might be able to help

Falconmaster:

https://www.mathway.com/BasicMath

kaylak:

i tried you have to pay for steps

Falconmaster:

if you have it slove it and show the steps

kaylak:

f(x)=square root 2x-6 find inverse

Falconmaster:

sorry thats all i can do i wiah i could help more

ThisGirlPretty:

@Vocaloid

kaylak:

@Layla help?

Layla:

hold on for 1 sec

kaylak:

gotta have steps

Layla:

okay

Vocaloid:

I'm on my phone so this might be hard to explain but You have to calculate f(g(x)) and g(f(x)) and if they both equal to x they are inverses

kaylak:

[ 5 [ x / [ 5x + 1] ] + 1 ] / [ x / [5x + 1] ] = [ (5x + 5x + 1 ) / [ 5x + 1) ] / [x / [5x + 1] ] = ( [ 10x + 1] / x ] = 10 + 1/x is this right

Layla:

yes

kaylak:

next question give me a sec

kaylak:

find the inverse function for squre root 2x-6

Vocaloid:

Let y = sqrt(2x-6) Switch y and x and then re-solve for y

kaylak:

To find the inverse interchange the variables and solve for y. y=square root 2x-6 x=square root 2y-6 divide each side by 2 f^-1(x)=x^2/2 +3 i feel like i missed a step

Vocaloid:

Why are you dividing by 2?

kaylak:

6/2=3 right

Vocaloid:

If you have x = sqrt(2y-6) you should square both sides

kaylak:

oh ok

kaylak:

and the rest is right?

Vocaloid:

Right The answer is correct but the steps are incomplete

kaylak:

f(x)=(x-2)^3+8 i have to find inverse and show that this is one to one?

Vocaloid:

Same process for finding the inverse I actually forgot how to show one to one so I will look it up

kaylak:

lol i know how to graph it and determine but not algebraically

Vocaloid:

Ok to show one to one you have to show that f^1(f(x)) = x

Vocaloid:

f^(-1) sorry

kaylak:

do i just multiply f and x?

Vocaloid:

You take the function f(x) and let it be your x value in the function f^-1

Vocaloid:

So you would replace x with (x-2)^3+8 in the inverse

kaylak:

To find the inverse interchange the variables and solve for y. y=(x-2)^3+8 now what would i do and how would i square the other problem i got x=x^2 and that's about it

Vocaloid:

x = (y-2)^3 +8 Subtract 8 from both sides then take the cube root

Ultrilliam:

Welcome to QC! ^_^

kaylak:

so -8=(y-2)=x

Vocaloid:

Not quite cuberoot(x-8) = y - 2

kaylak:

so next do we cube it ? i guess

Vocaloid:

No just solve it for y

Vocaloid:

All you have to do is add 2 to both sides

kaylak:

2+cube root -8+x?

Vocaloid:

Don't change what's inside the root symbol just add 2 outside

Vocaloid:

y = cuberoot(x-8) + 2

kaylak:

ok so cube x-8 +2

Vocaloid:

Good just make sure to write the correct root symbol

kaylak:

mathway has the same answer in a different form

Vocaloid:

Yeah that's normal, you can write using a different form as long as it's the same function

kaylak:

now one to one part of it

Vocaloid:

@sillybilly123 are you free to take over? I have an exam coming up

Vocaloid:

You just need to show that f^-1(f(x)) = x

sillybilly123:

nah, mate!! you've found the inverse. the OP now needs to go through the definition of a function. i had a play with the original question, which was quite diffrent, on Desmos and the **inverse function** is a function if you slice and dice. the one you've done looks good, IMVHO, and so should be OK. your exams are more important. go study!

kaylak:

i wass about to say didn't i technically prove that it's one-to-one